OpenStudy (shiree19):

Explain how to write a function rule from the table below. Then write a function rule. Table: x: y: 2 1 4 0 6 -1 Please explain how to do this so I can do it on my own in the future. Thanks.

OpenStudy (imstuck):

The first thing to do is to find the slope of the line using the points they provided. The slope formula is the difference of the y points (y2 - y1) over the difference of the x points (x2 - x1), since slope is rise over run and rise is the y axis and run is the x axis. I chose the two points (2, 1) and (4, 0). The y2 - y1 would be 0 - 1, and the x2 - x1 would be 4 - 2. 0 - 1/4 - 2 = -1/2. That's the slope of the line. Now to get an equation or a function for this. Use the point-slope format to do this part, which is\[y-y _{1}=m(x-x _{1})\] where m is the slope you just found and y1 and x1 are coordinates from any ONE of the points given in the table. Heres how it looks using point (4, 0):\[y - 0=-\frac{ 1 }{ 2 }(x-4)\]Now distribute the -1/2 into the parenthesis and get \[y-0=-\frac{ 1 }{ 2 }x+2\] Move the zero over to the other side to set the whole thing to = y. Most times this number will not be a zero, and moving it will affect the other side of the equation, but since it is zero, we get a final function or linear equation of\[y=-\frac{ 1 }{ 2 }x+2\]with your slope as -1/2 and your y-intercept as a positive 2. Do you see that?

OpenStudy (shiree19):

yes. thanks.

OpenStudy (shiree19):

But, I thought that a function rule was like: f(x) = .......... how does that factor in with this problem?